# Three-coloring triangle-free graphs without long forbidden paths

### v1

@@ 1 Introduction
If $\mathcal{F}$ contains exactly one graph $F$, it is known that for each integer $k\geq 3$, the $k$-coloring problem restricted to $F$-free graphs remains NP-complete unless $F$ is a disjoint union of paths [9, 17, 21]. The case where $F$ is connected (i.e., $F$ is a path) received the most attention. Huang [18] proved that the 4-coloring problem for $P_{7}$-free graphs and the 5-coloring problem for $P_{6}$-free graphs are both NP-complete. On the other hand, Hoàng, Kamiński, Lozin, Sawada and Shu [15] proved that the $k$-coloring problem for $P_{5}$-free graphs can be solved in polynomial time for any fixed $k\geq 5$, whereas Chudnovsky, Spirkl and Zhong [7, 8] showed that the $4$-coloring problem for $P_{6}$-free graphs can be solved in polynomial time. By combining these results, one obtains a complete classification of the complexity of $k$-colouring $P_{t}$-free graphs for any fixed $k\geq 4$ and $t\geq 1$. For $k=3$, the strongest known result is due to Bonomo, Chudnovsky, Maceli, Schaudt, Stein and Zhong [2] who proved that the 3-coloring problem for $P_{7}$-free graphs can be solved in polynomial time.
@@
The girth of a graph $G$ is the length of its shortest cycle (or $\infty$ if $G$ has no cycles). Hence, graphs with girth at least $g$ correspond with $\{C_{3},C_{4},\ldots,C_{g-1}\}$-free graphs, which generalizes the $C_{3}$-free setting we considered so far. For any positive integer $t$, let $g(t)$ be the minimum positive integer such that there are only finitely many 4-vertex-critical $P_{t}$-free graphs with girth at least $g(t)$. The results from Golovach, Paulusma and Song [14] imply that for all $t\geq 15$, we have $g(t)\leq\frac{4}{9}(t+3)+1-c_{t}$, where $c_{t}=\frac{2}{9}$ if $t\equiv 4\pmod{9}$ or $t\equiv 6\pmod{9}$ and $c_{t}=0$ otherwise. There are only few values of $t$ for which the exact value of $g(t)$ is known. From the results mentioned in our previous discussion, it follows that $g(t)=3$ when $1\leq t\leq 6$ [6] and $g(7)=4$ [6, 1]. In Section 4 we present the third contribution of the current paper and construct an infinite family of 4-vertex-critical $\{4K_{2},C_{3}\}$-free graphs (note that these graphs are also $\{P_{11},C_{3}\}$-free). This is the first infinite family of 4-vertex-critical $\{P_{t},C_{3}\}$-free graphs for any value of $t$. Goedgebeur and Schaudt [12] showed that $g(11)\leq 5$ and so our construction settles $g(11)=5$.
@@ 2 Graph generation algorithm
This settles the claim from the introduction that $g(11)=5$. We remark that it is not possible to replace $P_{11}$ by $P_{10}$ in Corollary 4.2, since the set $\{v_{1},v_{9},v_{3k+5},v_{3},v_{11},v_{3k+7},v_{5},v_{13},v_{3k+9},v_{7}\}$ induces a $P_{10}$ in $G_{k}$ for each $k\geq 4$.
@@ 5 Conclusion
In this paper, we presented several contributions for the $3$-coloring problem on $\{F,C_{3}\}$-free graphs, where $F$ is an induced subgraph of a path. More precisely, we showed that there are exactly three 4-vertex-critical $\{P_{7},C_{3}\}$-free graphs containing an induced $C_{7}$, that every $\{P_{5}+P_{1},C_{3}\}$-free graph is $3$-colorable and that there is an infinite family of 4-vertex-critical $\{4K_{2},C_{3}\}$-free (and hence also $P_{11}$-free) graphs. Regarding our last result, we wish to point out that most infinite families of $k$-vertex-critical $\mathcal{F}$-free graphs from the literature are either circulant graphs or obtained by adding universal vertices to circulant graphs [3, 5, 6, 16]. Hence, all such graphs have girth at most $4$ when $k\geq 4$. Moreover, Golovach, Paulusma and Song [14] proved that $k$-coloring can be solved in polynomial time for $\{P_{t},C_{4}\}$-free graphs for any $t,k\geq 1$. These observations lead us to the following question that sparks our interest.
@@
Are there integers $t,k\geq 1$ such that there are infinitely many $k$-vertex-critical $\{P_{t},C_{3},C_{4}\}$-free graphs?
@@ Acknowledgements
We acknowledge the support the joint FWO-NSFC scientific mobility project with grant number VS01224N. The research of Jan Goedgebeur was supported by Internal Funds of KU Leuven and an FWO grant with grant number G0AGX24N. Jorik Jooken is supported by a Postdoctoral Fellowship of the Research Foundation Flanders (FWO) with grant number 1222524N.

### v2

@@ Abstract
Keywords: Graph coloring; Hereditary graph classes; Vertex-critical graphs
@@ 1 Introduction
If $\mathcal{F}$ contains exactly one graph $F$, it is known that for each integer $k\geq 3$, the $k$-coloring problem restricted to $F$-free graphs remains NP-complete unless $F$ is a disjoint union of paths [8, 14, 19]. The case where $F$ is connected (i.e., $F$ is a path) received the most attention. Huang [15] proved that the 4-coloring problem for $P_{7}$-free graphs and the 5-coloring problem for $P_{6}$-free graphs are both NP-complete. On the other hand, Hoàng, Kamiński, Lozin, Sawada and Shu [13] proved that the $k$-coloring problem for $P_{5}$-free graphs can be solved in polynomial time for any fixed $k\geq 5$, whereas Chudnovsky, Spirkl and Zhong [6, 7] showed that the $4$-coloring problem for $P_{6}$-free graphs can be solved in polynomial time. By combining these results, one obtains a complete classification of the complexity of $k$-coloring $P_{t}$-free graphs for any fixed $k\geq 4$ and $t\geq 1$. For $k=3$, the strongest known result is due to Bonomo, Chudnovsky, Maceli, Schaudt, Stein and Zhong [2] who proved that the 3-coloring problem for $P_{7}$-free graphs can be solved in polynomial time.
@@
The girth of a graph $G$ is the length of its shortest cycle (or $\infty$ if $G$ has no cycles). Hence, graphs with girth at least $g$ correspond with $\{C_{3},C_{4},\ldots,C_{g-1}\}$-free graphs, which generalizes the $C_{3}$-free setting we considered so far. For any positive integer $t$, let $g(t)$ be the minimum positive integer such that there are only finitely many 4-vertex-critical $P_{t}$-free graphs with girth at least $g(t)$. The results from Kamiński and Pstrucha [17] imply that for all $t\geq 1$, we have $g(t)\leq 5$. There are only few values of $t$ for which the exact value of $g(t)$ is known. From the results mentioned in our previous discussion, it follows that $g(t)=3$ when $1\leq t\leq 6$ [5] and $g(7)=4$ [5, 1]. In Section 4 we present the third contribution of the current paper and construct an infinite family of 4-vertex-critical $\{4K_{2},C_{3}\}$-free graphs (note that these graphs are also $\{P_{11},C_{3}\}$-free). This is the first infinite family of 4-vertex-critical $\{P_{t},C_{3}\}$-free graphs for any value of $t$. Hence, by combining our result with the result by Kamiński and Pstrucha [17], we show that $g(t)=5$ for all $t\geq 11$.
@@ 2 Graph generation algorithm
This settles the claim from the introduction that $g(t)=5$ for all $t\geq 11$. We remark that it is not possible to replace $P_{11}$ by $P_{10}$ in Corollary 4.2, since the set $\{v_{1},v_{9},v_{3k+5},v_{3},v_{11},v_{3k+7},v_{5},v_{13},v_{3k+9},v_{7}\}$ induces a $P_{10}$ in $G_{k}$ for each $k\geq 4$.
@@ 5 Conclusion
In this paper, we presented several contributions for the $3$-coloring problem on $\{F,C_{3}\}$-free graphs, where $F$ is an induced subgraph of a path. More precisely, we showed that there are exactly three 4-vertex-critical $\{P_{7},C_{3}\}$-free graphs containing an induced $C_{7}$, that every $\{P_{5}+P_{1},C_{3}\}$-free graph is $3$-colorable and that there is an infinite family of 4-vertex-critical $\{4K_{2},C_{3}\}$-free (and hence also $P_{11}$-free) graphs. The following natural question sparks our interest:
@@
What is the largest integer $t$ for which there are only finitely many 4-vertex-critical $\{P_{t},C_{3}\}$-free graphs?
@@ Acknowledgements
We acknowledge the support the joint FWO-NSFC scientific mobility project with grant number VS01224N. The research of Jan Goedgebeur was supported by Internal Funds of KU Leuven and an FWO grant with grant number G0AGX24N. Jorik Jooken is supported by a Postdoctoral Fellowship of the Research Foundation Flanders (FWO) with grant number 1222524N. We are grateful to Ben Cameron for making us aware of the result by Kamiński and Pstrucha [17] and for suggesting Question 1 in the last section.